OpenStudy (anonymous):

divide: (x^2-6x-7)/(3x^2-147) divided by (x^2+8x+7)/(x^2+14x+49) here's what i have so far-- (x^2-6x-7)/(3x^2-147) * (x^2+14x+49)/(x^2+8x+7) i factored it out-- (x+1)(x-7)/3x^2-147 * (x+7)(x+7)/(x+1)(x+7) what im getting stuck on is the 3x^2-147. I know i have to get it to 3(x^2-27), but i cant factor it further after that, right? and my online homework program wont accept it as just 3(x^2-27). I tried to find factors of 27 but it's only 1, 3, 9 and 27. Am I missing something completely obvious? I'm at a loss here.

OpenStudy (anonymous):

you sure it is a square and not a cube?

OpenStudy (anonymous):

\[x^3-27\] factors because it is the difference of two cubes just asking, maybe it is not

OpenStudy (anonymous):

it's definitely a square

OpenStudy (anonymous):

yes, so it is

OpenStudy (anonymous):

hmm apparently the correct answer is \(\frac{1}{3}\) so lets see how we can get that ...

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

oh gosh okay

OpenStudy (anonymous):

it is not \[3x^2-147=3(x^2-27)\] it is \[3x^2-147=3(x^2-49)\]

OpenStudy (anonymous):

arithmetic will screw you up every time you good from there?

OpenStudy (anonymous):

.. yes omg,, of course it was a simple mistake wow thank you for your help